OpenStudy (anonymous):

Solve. 3/x+5/x=-4 x=-1/2 x=2 x=1/2 x=-2

OpenStudy (anonymous):

hi again

OpenStudy (anonymous):

\[\frac{3}{x}+\frac{5}{x}=\frac{8}{x}\]

OpenStudy (anonymous):

i think i wrote the question wrong

OpenStudy (anonymous):

\[\frac{8}{x}=-4\iff -4x=8\iff x=-2\]